# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

### v2

@@ 6.3 Comparison of Harmful Responses
In this section, we conduct a qualitative analysis of our attack’s outputs and how they differ from the original model and baseline methods. Table IV provides examples that highlight the general trends we observe for each method.
Our attack generates higher-fidelity responses than Baseline BFA. PrisonBreak induces a jailbroken state largely avoidant of the failure modes presented in §5.1. Models rarely backtrack affirmative responses, and although we find a few cases of overfitting, they are still accompanied by harmful content. Conversely, Baseline BFA is severely susceptible to overfitting and commonly yields no harmful content but rather the repetition of a few phrases. Despite these undesirable artifacts, the HarmBench evaluator tends to still report these instances as harmful, likely because the model does not explicitly deny the query and repeatedly affirms it. As a result, the ASRs reported for Baseline BFA likely overestimate the amount of harmful content produced.
Comparison with non-BFA methods. While all three attacks (excluding Baseline BFA) often generate high-fidelity content, they differ with respect to output style. GCG frequently involves portions of the adversarial suffix, e.g., the model formats the example response in “Irish colloquial style” (the suffix contains “Irish”). Models sometimes overvalue the suffix and ignore the harmful request completely, possibly playing a role in the prompt-based method’s lower ASR. The responses from PrisonBreak are semantically closest to Ortho; models attacked by these methods provide very similar step-by-step instructions and generally respond to queries with comparable ideas.
@@ 6.4 Exploitability Under Black-Box Settings
We also evaluate the effectiveness in black-box scenarios where an adversary may not have white-box access to the exact model they wish to jailbreak. Here, the attacker can exploit the standard paradigm of LLM development: pretraining-then-finetuning. Because pre-training is the primary computational bottleneck, many practitioners open-source pre-trained LLMs for others to fine-tune [30, 25, 29, 59, 31, 60]. This practice has led to a proliferation of models that share components, such as architectures, pre-training, fine-tuning, or post-training alignment procedures. As a result, an attacker with white-box access to an open-source model (source) can identify effective bit-flip locations in memory and transfer them to a target model in a black-box setting without any access to its parameters.
Table V summarizes our results. We first select source-target model pairs from our experiments in §6.2. For each pair, we induce bit-flips to a target model at the locations identified from the corresponding source model. To further isolate the impact of post-training alignment, we also include the SFT-only checkpoint of Tulu3-8b, Tulu3-8b-SFT, to assess effectiveness when all factors except post-training alignment are held constant. We discover that bit-flip locations transfer effectively across models that share key stages of the LLM development pipeline. In the case of Tulu3-8b-SFTTulu3-8b, where the only difference between the models is post-training alignment, the attack achieves the highest transferability with an ASR of 70.4%. This is comparable to Baseline BFA and reaches 81% of the white-box attack’s effectiveness. Vicuna-7b-v1.5Llama2-7b, where the models share both architecture and pre-training, yields an ASR of 49.1%, outperforming Baseline BFA and achieving 56% of the white-box performance. This result is particularly alarming: since a common practice in LLM development is to fine-tune or safety-align pre-trained models—as seen in Vicuna, Alpaca, and Tulu, which are all fine-tuned variants of Llama—our attack is effective even in black-box scenarios.
However, as expected, bit-flip locations do not transfer well when the source and target models differ substantially in their LLM development procedures. Vicuna-7b-v1.3Vicuna-7b-v1.5 yields an ASR of 19.5%, as the two models share the same architecture but differ in pre-training, fine-tuning, and post-training alignment. Qwen2-1.5bQwen2-7b results in an ASR of 16.4%, likely due to architectural differences. In both cases, differences in architecture or development procedures likely lead to substantial divergence in learned parameters, thereby reducing the effectiveness of transferred bit-flips.
@@ 6.5 Impact of Attack Configurations
We analyze how sensitive our ASR is to different attack configurations. We show our results on Llama2-7B in Figure 2. Please refer to Appendix B for another model.
Impact of the proxy dataset size. Our default proxy dataset size is 41—a small number compared to prior work which uses 50–2500 samples to compute the attack objective. The leftmost figure shows the ASR and Acc. for different proxy dataset sizes in jailbreaking Llama2-7B. We observe similar output quality across all dataset sizes and find that any dataset size over 5 achieves an ASR above 85%, indicating that a large dataset is not necessary for successful jailbreaks. However, we find that smaller datasets are often dependent on the specific samples they contain; we use the full proxy dataset to promote consistent results between models.
Impact of , the number of critical weights selected per layer during the search process. A smaller reduces the total computations, but may compromise attack success and model utility. The second figure from the left shows the impact of on ASR and Acc. As increases from 10 to 150, ASR increases from 73.6% to 95.0%. But, Acc. decreases marginally from 51.8% to 50.7%, and the computation required for our attack increases proportionally. Despite the reduction in Acc., the quality of harmful responses remains similar across . Our preliminary investigation finds that 50–100 balances search time while increasing ASR only with a marginal reduction of Acc. We use .
Impact of , the utility score threshold used when selecting target bits. We vary in , with smaller values indicating that flipping the selected bit only slightly decreases model utility. The third figure from the left shows the results: does not affect ASR, but as it is increased Acc. is reduced from 52.9% to 50.3%. Manual analysis of outputs across values shows that a smaller threshold produces higher-fidelity responses for both benign and harmful queries. For example, the generations of Llama2-7b with contain much more overfitting than those with lower thresholds, even at . In our attacks, we set .
@@
7 Rowhammer Exploitability
We now demonstrate the Rowhammer exploitability of PrisonBreak. Rowhammer exploitation typically breaks down into three critical phases: (1) templating, (2) memory massaging, and (3) reflipping. Templating identifies DRAM rows and bit positions that reliably flip under repeated access (hammering). Attackers test different aggressor rows using various access patterns to discover potentially exploitable bit positions [39]. Prior studies show that templating takes anywhere from several hours to days [1, 39] but is performed offline and only once per machine. Once exploitable bit positions are identified, memory massaging relocates the pages containing the target bits in a model into physical DRAM rows containing these vulnerable bits. This is achieved through direct memory control and hammering [36], or by repeatedly allocating and freeing “dummy” pages to force the target into the desired location [61, 62]. With the target page in place, the attacker reflips the vulnerable bit to corrupt model parameters. This final phase is performed online and typically completes in a few minutes [1, 63, 12]. Attacks are carefully crafted to hammer only templated aggressors [62, 64, 65], so as to avoid extraneous flips that could crash the process or degrade model utility.
Prior work has shown that precise Rowhammer exploitation is feasible for neural networks [1, 12, 7, 11], even when flipping up to 59 more bits [11] than PrisonBreak requires. Rather than re-demonstrating Rowhammer on limited commodity hardware, we follow recent work [66] and use simulations to assess the exploitability across dozens of DRAM modules. This prevents our results from being an one-time exploitation and instead allows us to comprehensively assess generalizability to many DRAM modules, in line with prior large-scale Rowhammer studies [39].
Models # of Bit-Flips DDR4 LPDDR4 High-RV Mid-RV High-RV A17 (65) A8 (43) A12 (20) A19 (5.4) D1 (3.1) A11 (1.5) 15/16 Modules (19.5) C5 (10.8) Vicuna-7B-v1.3 5 100% 100% 100% 99.6% 96.3% 38.1% 100% 100% Vicuna-7B-v1.5 12 100% 100% 100% 93.5% 59.8% 0% 100% 100% Vicuna-13B-v1.5 20 100% 100% 100% 97% 76% 0.4% 100% 98% Llama2-7B 25 100% 100% 100% 19.4% 0.4% 0% 100% 80% Llama2-13B 18 100% 100% 100% 96.9% 74.8% 0.1% 100% 100% Llama3-8B 20 100% 100% 100% 94.7% 63.2% 0.1% 100% 100% Llama3-70B 18 100% 100% 100% 95.1% 68.2% 0.1% 100% 100% Tulu3-8B 13 100% 100% 100% 95.4% 66.3% 0% 100% 99% Qwen2-1.5B 25 100% 100% 100% 93.8% 53.5% 0% 100% 100% Qwen2-7B 25 100% 100% 100% 93.5% 56.1% 0% 100% 100%
7.1 Experimental Setup and Methodology
To test the exploitability, we assume the previously demonstrated templating, massaging, and re-flipping capabilities [1, 7, 12] and simulate an attacker’s ability to locate and re-flip vulnerable bit locations identified by PrisonBreak. Specifically, we conduct Monte Carlo simulations based on Rowhammer profiles from Blacksmith [39]. We evaluate 40 DDR4 and 16 LPDDR4X (Low-Power DDR4 eXtension) DRAM modules from four manufacturers. Each DRAM module consists of a profile detailing the number of cells vulnerable to Rowhammer and the number of bit-flips possible in each flip direction (01 or 10). The total number of vulnerable cells determines a module’s Rowhammer vulnerability (RV), defined as the ratio of cells susceptible to bit-flips to the total number of DRAM cells. We categorize the 40 DDR4 modules into three RV levels:
High: RV 10 parts per million (ppm).
Mid: 1 RV 10 ppm.
Low: RV 1 ppm.
High-RV modules show vulnerability levels comparable to those exploited in prior BFA studies [1, 12, 11]; Mid- and Low-RV modules have 3–14 and 18–2560 fewer vulnerable cells than the High-RV modules. All 16 LPDDR4X modules have High-RVs. Our simulation considers a system with a single-channel configuration for all modules. Descriptions of these modules are available in prior work [39].
Methodology. In each iteration of the Monte Carlo simulation, we randomly initialize the Rowhammer cells based on the RV of each DRAM module. Cell locations are sampled uniformly, because prior work [1] found that flippable bit-flip locations are uniformly distributed, and the Rowhammer profiles we reference [39] do not specify the exact locations of vulnerable cells. After the cell initialization, we check if there are any exploitable Rowhammer cells for our target bits, following a process similar to the Hammertime framework [63]. Each target bit must be matched to a cell with the same offset and flip direction, otherwise it is not considered flippable. This constraint mimics the attacker’s memory massaging step, where they can only massage target bits into specific DRAM cells. The attack is successful if we find exploitable locations for all required target bits necessary to jailbreak the LLM, modeling a successful memory massaging attempt by the attacker that can be followed by Rowhammer. We run the simulations for 100,000 iterations.
Metrics. We employ two metrics to quantify the Rowhammer exploitation. The first is the Rowhammer success rate (RSR), estimated using our Monte Carlo simulation. RSR is defined as the ratio of successful trials to the total number of trials. This metric applies to the hardware-agnostic attacker, who does not profile the DRAM before attacking, making their RSR dependent on the RH vulnerability of the DRAM module. The second metric is the attack success rate (ASR) introduced in §6.1. We use this metric for the hardware-aware attacker; this attacker always selects flippable target bits, making Monte Carlo simulations unnecessary.
7.2 Evaluation Results
Hardware-agnostic attacker. We first evaluate the scenario where the attacker identifies a set of target bits and aims to flip them across multiple DRAM modules. We use the same target bits as in §6.2. The third and fourth columns in Table VI show the RSR for three representative DDR4 modules at the High- and Mid-RV levels; we omit the Low-RV results due to low RSR (–%) and present them with the full results in Table X in the Appendix. Our hardware-agnostic attack achieves an RSR of 100% across all High-RV DRAMs, in alignment with prior work [1, 12, 11, 7], especially as our attack requires a similar or fewer number of bit-flips. The attack is reasonably successful on Mid-RV DRAMs, with an RSR of 96.3% for Vicuna-7B-v1.3 on D1, a module with 5.4 less RV than those used in prior work. Even for Qwen2-7B, which requires 25 bit-flips, we achieve an RSR of 56.1%. But, we are less successful for Llama2-7B against Mid-RV modules as many target bits share the same offsets, requiring multiple Rowhammer cells with identical offsets within a DRAM row. Given the uniform distribution of vulnerable Rowhammer cells, this lowers the probability of finding exploitable locations. The attacks are not successful on Low-RV DRAMs, as these modules are significantly more resilient to Rowhammer with far fewer exploitable bits (at most several thousand).
The rightmost column in Table VI shows our results on LPDDR4X modules. Our hardware-agnostic attack is consistently successful with an RSR of 80–100% across all settings. This is because LP devices are more vulnerable to Rowhammer due to power and area constraints [39, 37].
Modules Vicuna-7B-v1.5 Llama3-8B Qwen2-7B C4 (0.4) 94.34 82.39 88.68 A11 (1.5) 90.56 78.62 90.57
Hardware-aware attacker. This attacker identifies target bits that map to vulnerable cells and can succeed even against secure DRAM modules. We evaluate against the A11 (Mid-RV) and C4 (Low-RV) modules, where the hardware-agnostic attack achieved very low RSR (0–38.1%). Table VII shows the ASR for the hardware-aware attacker, computed on HarmBench as in §6. We do not report the RSR as this attack produces bit-flip sequences that are always flippable in the target DRAM. The hardware-aware attacker achieves a high ASR on both modules, comparable to the upper bound results we obtain in §6. Even in constrained settings, PrisonBreak can still jailbreak models with up to 25 bit-flips. For reference, the A11 and C4 modules are 12–46 more secure than those used in prior work [1].
@@ 8.1 Structural Analysis
Here we analyze bit-flips that led to the highest reduction in our jailbreak objective,
particularly the top 1% of bits ranked by the jailbreak score. We consider all 16 bit locations in each critical weight. Figure 3 shows the proportion of target bits across different components of Llama2-7B.
Bit locations. The leftmost figure confirms that large parameter changes, induced by flipping the three most significant bits (MSBs) in the exponent, lead to the greatest reduction in our objective. Within the top 1% bits across each bit location in the half-precision floating-point weight value, we see that the three most significant exponent bits makeup over 94% of all top bits flips across iterations. Of these three bit locations, the MSB in the exponent is likely chosen (with a proportion of 48.0%) in the first round, while the second bit is more likely for the following iterations with proportions of 42.4–60.4%. This is because, while flipping the second bit induces smaller parameter changes than the first, it prevents excessive changes that often reduce a model’s generation capabilities. Our findings corroborate the actual bits selected: 31.5% are MSBs, and 64% are the second exponent bits.
Types of layers composing the Transformer block. Our attack identifies more candidate bits in certain Transformer layers than others. The middle figure shows the proportion of the top 1% bits found in each layer type. Overall, the linear layer for the value in multi-head attention (attn.v) has relatively higher proportions (25.2–35.0%); notably, 28.5% of all bits selected in §6 come from these layers. This may be because PrisonBreak primarily affects how the model responds to harmful queries (i.e., not refusing to answer) rather than how it recognizes them. In multi-head attention, “how to respond” corresponds to the value layers, while “how to recognize” aligns with the key and query layers.
Layer positions. Prior work [67] shows that in Transformer-based models, earlier layers handle basic concepts like word order, middle layers focus on syntax, and final layers are task-specific. Building on this, we analyze the relationship between layer positions and the likelihood of selecting target bits in Llama2-7B. We categorize layers into early (blocks 1–10), middle (blocks 11–22), and final (blocks 23–32). The rightmost figure shows the results. Early layers are chosen less often, accounting for under 25% of top bits in most iterations. Middle and final layers dominate, particularly with middle layers contributing 33.5–63.1%. Our findings corroborate the prior work: in early iterations, the attack is more likely to select bits from the final layers, as it focuses on the task of generating harmful responses. In the later iterations, the attack shifts to choosing target bits from the middle layers, focusing on generating high-quality responses while maintaining attack success (the attack objective).

### v3

@@ 6.3 Comparison of Harmful Responses
We conduct a qualitative analysis of our attack’s outputs and how they differ from the original model and baseline methods. Table 4 provides examples that highlight the general trends.
Our attack generates higher-fidelity responses than Baseline BFA. PrisonBreak induces a jailbroken state largely avoidant of the failure modes presented in §5.1. Models rarely backtrack affirmative responses, and although we find a few cases of overfitting, they are still accompanied by harmful content. Conversely, Baseline BFA is severely susceptible to overfitting and commonly yields no harmful content but rather nonsensical repetition. Despite this, the HarmBench evaluator tends to still report many instances as harmful, likely because the model does not explicitly deny the query. As a result, the ASRs reported for Baseline BFA likely overestimate the true amount of harmful content produced.
Comparison with non-BFA methods. While all three attacks (excluding Baseline BFA) often generate high-fidelity content, their output style differs. GCG frequently involves portions of the adversarial suffix, e.g., the model formats the example response in “Irish colloquial style” (the suffix contains “Irish”). Models sometimes overvalue the suffix and ignore the harmful request entirely, possibly contributing to the method’s lower ASR. The responses from PrisonBreak are semantically closest to Ortho; models attacked by these methods provide very similar step-by-step instructions and generally respond to queries with comparable ideas.
@@ 6.4 Exploitability Under Black-Box Settings
We evaluate the effectiveness in black-box scenarios where an adversary may not have white-box access to the exact model they wish to jailbreak. Here, the attacker can exploit the standard paradigm of LLM development: pretraining-then-finetuning. Because pre-training is the primary computational bottleneck, many practitioners open-source pre-trained LLMs for others to fine-tune [25, 81, 82, 8, 92, 79]. This has led to a proliferation of models that share components, such as architectures, pre-training, fine-tuning, or alignment procedures. Thus, an attacker with white-box access to an open-source model (source) can identify effective bit-flip locations in memory and transfer them to a target model in a black-box setting without any access to its parameters.
Table 5 summarizes our results. We select source–target model pairs from §6.2. For each pair, we flip bits in the target model at locations identified from the source model. To isolate the impact of post-training alignment, we also include the SFT-only checkpoint of Tulu3-8b (Tulu3-8b-SFT) to assess effectiveness when all factors but post-training alignment are held constant. We find that bit-flip locations transfer effectively across models that share key stages of the LLM development pipeline. For Tulu3-8b-SFTTulu3-8b, where the only difference is post-training alignment, we achieve the highest transferability with an ASR of 70.4%. This is comparable to Baseline BFA and reaches 81% of the white-box attack’s effectiveness. Vicuna-7b-v1.5Llama2-7b, where the models share both architecture and pre-training, yields an ASR of 49.1%, outperforming Baseline BFA and reaching 56% of the white-box performance. These results show that the practice of fine-tuning or aligning widely available pre-trained models (e.g., Vicuna, Alpaca, and Tulu, are all fine-tuned variants of Llama) makes our attack effective even in black-box scenarios.
However, bit-flip locations do not transfer well when the source and target models differ substantially in their development procedures. Vicuna-7b-v1.3Vicuna-7b-v1.5 yields an ASR of 19.5%, as the two models share the same architecture but differ in pre-training, fine-tuning, and post-training alignment. Qwen2-1.5bQwen2-7b results in an ASR of 16.4%, likely due to architectural differences. In both cases, differences in architecture or development procedures result in a substantial divergence in learned parameters, thereby reducing the effectiveness of transferred bit-flips.
@@ 6.5 Impact of Attack Configurations
We analyze how sensitive our ASR is to different attack configurations. We show our results on Llama2-7B in Figure 2. Please refer to Appendix F for another model; we also study the impact of model datatype in Appendix E.
Impact of the proxy dataset size. Our default proxy dataset size is 41—a small number compared to the 50–2500 samples used in prior work. The leftmost figure shows the ASR and Acc. for different proxy dataset sizes in jailbreaking Llama2-7B. We observe similar output quality across all sizes and find that any dataset size over 5 achieves an ASR above 85%, indicating that a large dataset is not necessary for successful jailbreaks. However, we find that smaller datasets are often dependent on the specific samples they contain; we use the full proxy dataset to promote consistent results between models.
Impact of , the number of critical weights examined per layer. A smaller reduces the total computations, but may compromise ASR. The second figure from the left shows the impact of on ASR and Acc. As increases from 10 to 150, ASR increases from 73.6% to 95.0%. But, Acc. decreases marginally from 51.8% to 50.7%, and the required compute increases proportionally. Despite the reduction in Acc., the quality of harmful responses remains similar across . Our preliminary investigation finds that 10–100 yields strong results; we conservatively use 100 for most experiments.
Impact of , the utility score threshold used when selecting target bits. We vary in , with smaller values indicating less impact on model utility. The third figure from the left shows the results: does not affect ASR, but as it is increased Acc. is reduced from 52.9% to 50.3%. Manual analysis of outputs across values shows that a smaller threshold produces higher-fidelity responses for both benign and harmful queries. For example, the generations of Llama2-7b with contain much more overfitting than those with lower thresholds, even at . In our attacks, we set .
@@
7 End-to-End Exploitation
In this section, we demonstrate the practicality of PrisonBreak in an end-to-end setting: we jailbreak language models using Rowhammer bit-flips on an NVIDIA A6000 GPU with 48GB GDDR6 DRAM. While Lin et al. [51] present a proof-of-concept exploit that flips random bits to degrade model utility, to our knowledge, this is the first targeted Rowhammer-based bit-flip attack conducted on a GPU.
7.1 Attack Setup
Following our threat model in §3, we assume an attacker is co-located with a victim language model on a time-sliced GPU, where RMM [1] manages memory allocations and memory swapping [23] offloads idle data when GPU memory is full. The attack proceeds in two phases. In the offline phase—where co-location with the victim is not required—we profile the GPU DRAM and apply our hardware-aware search to identify target bits. In the subsequent online phase, we exploit Rowhammer to flip these bits, thereby directly tampering with the victim’s model. We detail each step below:
DRAM profiling. To identify vulnerable bit locations in the GPU, the attacker launches a hammering campaign. Following the procedure of Lin et al. [51], we use DRAMA [64] to reverse-engineer row addressing, multi-warp hammering to increase activation counts, and synchronized, single-sided hammering patterns to bypass in-DRAM TRR. The resulting profile records the offset and flip direction of each bit-flip, along with the aggressor pattern that induces it.
An additional constraint for the exploit is that the attacker must be able to access the row neighboring the target victim row. RMM allocates each model layer for the victim as a contiguous block in memory, and prior work [51] shows that neighboring rows map to addresses within 2MB. With single-sided hammering, in which aggressors lie on only one side of a target bit, this means that each bit-flip can only target weights within the first or last 2MB of a victim layer. To account for this, we record for each bit-flip the side (front or back) of the aggressor row and its distance from the target. This information enables our search process to determine which flips we can use to hammer which weights.
Target bit search. Due to the limited number of bit-flips on GDDR6 GPUs [51], we use our hardware-aware search algorithm to locate target bits. As described in §5.3, the attacker restricts their targets to bits whose offset and flip direction are in the DRAM profile. To ensure aggressors remain accessible for hammering, the attacker further limits target weights to the beginning (for front aggressors) or end (for back aggressors) of each layer, with the maximum allowable distance given by the row–aggressor distance of each profiled bit-flip.
Online exploitation. Finally, the attacker uses Rowhammer to flip target bits and jailbreak the LLM. This step presents two challenges. First, vulnerable bit-flips occur at fixed physical locations, while the victim’s model is mapped arbitrarily in memory; without placement control, victim weights may not align with vulnerable bits. Second, unlike prior demonstrations on CPU DRAM [94, 68, 80, 45, 50], GDDR6 DRAM has far fewer exploitable locations. To address these challenges, we employ a three-step online exploitation that enables precise hammering and physical bit-flip reuse:
Memory massaging. To flip target bits, the attacker must first massage the victim’s data into vulnerable locations. Following prior work [51], we exploit RMM’s immediate reuse of freed memory. To prepare a target bit for flipping, the attacker allocates large contiguous chunks of memory and then frees the chunk containing the desired flip. When the victim requests memory to load their model, RMM allocates the recently freed chunk first, thereby mapping the victim’s data to an attacker-chosen physical location.
Hammering. With the target bit in the desired physical location, the attacker uses the corresponding aggressor pattern from the DRAM profile to flip it with Rowhammer.
Memory swapping. We exploit GPU memory swapping to reuse the same physical bit-flip across multiple target weights. After hammering a target bit, the attacker allocates a large memory chunk to trigger swapping. When the victim’s process idles between time slices, the GPU offloads the model to the CPU, freeing all associated memory chunks. This restores access to the entire pool of physical bit-flips, which can be re-massaged via .
We repeat steps – for each target bit to jailbreak the victim model at runtime until a complete reload is performed.
7.2 Attack Effectiveness
We launch our hammering campaign on the A6000 to profile the DRAM, targeting eight banks. In total, we identify ten vulnerable bit locations (details provided in Appendix B). Compared to prior attacks on CPU DRAM [94, 68, 80, 45, 50], our target GPU DRAM exhibits almost – fewer bit flips, making our search space far more constrained. After profiling, we apply PrisonBreak’s target bit search to the victim model; since very few bit-flips are possible, we use to ensure a sufficient candidate set. We then perform the online exploitation to flip each identified bit via Rowhammer and our memory massaging techniques. After flipping up to 25 vulnerable bits, we compute the ASR and Acc. for each model using the evaluation procedure described in §6.
| Models | ASR | Acc. | P | #B |
| Vicuna-7B-v1.3 | 91.19 | -2.47 | 19 | 2 |
| Llama2-13B | 77.36 | -2.80 | 20 | 2 |
| Llama3-8B | 70.44 | -6.04 | 25 | 2 |
| Tulu3-8B | 81.13 | -2.74 | 20 | 2 |
| Qwen2-7B | 69.18 | -2.42 | 24 | 2 |
Results. Table 6 presents our results for one model from each family evaluated in §6. All bit-flips are successfully induced through hammering, confirming that our hardware-aware search precisely identifies only flippable bit-locations. PrisonBreak reliably jailbreaks LLMs in the end-to-end attack setting. Across all models, we achieve an ASR of 69.2–91.2%, which is comparable to the 80.5–98.1% obtained via our hardware-agnostic search. Notably, PrisonBreak attains these high ASRs using only two physical bit locations (#B). Although we identify ten vulnerable flips in profiling, the offline search process consistently selects only the two that map to the most significant exponent bit with flip directions. This demonstrates that PrisonBreak remains effective even against highly secure DRAM.
The impact on utility is modest, with an average Acc. of -3.3%, slightly larger than the -2.3% observed with the hardware-agnostic search. This difference likely stems from requiring a few additional bit-flips to reach high ASR (average P increases from 18 to 22), which introduces more collateral damage. Nevertheless, the drop is minor (within 1%), and we still preserve utility 4.7 better than Baseline BFA.
@@ 8.1 Structural Analysis
Here, we analyze bit-flips that lead to the highest reduction in our jailbreak objective,
particularly the top 1% of bits ranked by the jailbreak score (Def. 5.1). We consider all 16 bit locations in each critical weight. Figure 3 shows the proportion of target bits across different components of Llama2-7B.
Bit locations. The leftmost figure confirms that large parameter changes, induced by flipping the three most significant bits (MSBs) in the exponent, lead to the greatest reduction in our objective. Within the top 1% bits across each bit location in the float16 weight value, we see that the three most significant exponent bits make up over 94% of all top bits-flips across iterations. Of these three bits, the MSB in the exponent is likely chosen (with a proportion of 48.0%) in the first round, while the second bit is more likely for the following iterations with proportions of 42.4–60.4%. This is because, while flipping the second bit induces smaller parameter changes than the first, it prevents excessive damage to a model’s generation capabilities. Our findings corroborate the actual bits selected in §6: 31.5% are MSBs, and 64% are the second MSB.
Types of layers composing the Transformer block. Our attack identifies more candidate bits in certain layers. The middle figure shows the proportion of the top 1% bits found in each layer type. The linear layer for the value in multi-head attention (attn.v) has relatively higher proportions (25.2–35.0%); notably, 28.5% of all bits selected in §6 come from these layers. This may be because PrisonBreak primarily affects how the model responds to harmful queries (i.e., not refusing) rather than how it recognizes them. In multi-head attention, “how to respond” corresponds to the value layers, while “how to recognize” aligns with the key and query layers.
Layer positions. Prior work [72] shows that in Transformer-based models, earlier layers handle basic concepts like word order, middle layers focus on syntax, and final layers are task-specific. Building on this, we analyze the relationship between layer positions and the likelihood of selecting target bits in Llama2-7B. We categorize layers into early (blocks 1–10), middle (blocks 11–22), and final (blocks 23–32). The rightmost figure shows the results. Early layers are chosen less often, accounting for under 25% of top bits in most iterations. Middle and final layers dominate, particularly with middle layers contributing 33.5–63.1%. Our findings corroborate prior work: in early iterations, the attack is more likely to select bits from the final layers, as it focuses on the task of generating harmful responses. In later iterations, it shifts to choosing target bits from middle layers, focusing on generating high-quality responses while maintaining attack success.